# Forget What You Know about LLMs Evaluations - LLMs are Like a Chameleon

### v1

@@ 4.2 Models Under Evaluation
Table 1 provides an overview of the LLMs evaluated in our experiments. Our study covers a diverse set of architectures and parameter scales ranging from 1B to 236B parameters. This broad selection enables an in-depth analysis of how both architectural choices and model scale affect robustness to prompt perturbations.
Table 1: Overview of the evaluated LLMs. Models are grouped by family, model version, and the number of parameters (in billions).
Family Version Params
Qwen Qwen2.5 1.5B Yang et al. (2024) 1.5
Qwen2.5 3B 3
Qwen2.5 7B 7
Qwen2.5 32B 32
Qwen2.5 72B 72
Llama 3 Llama 3.2 1B Dubey et al. (2024) 1
Llama 3.2 3B 3
Llama 3.1 8B 8
Gemma Gemma 2 2B Team et al. (2024) 2
Gemma 7B 7
Gemma 27B 27
Phi Phi 3.5 4B Abdin et al. (2024) 4
Phi 4 15B 15
DeepSeek DeepSeek 7B Bi et al. (2024) 7
DeepSeek V2 16B 16
DeepSeek 236B 236
Yi Yi 6B Young et al. (2024) 6
Yi 9B 9
Others Apollo2 7B Zhu et al. (2024b) 7
Aquila 7B Zhang et al. (2024b) 7
Bloomz 7B Zhu (2023) 7
Falcon 7B Almazrouei et al. (2023) 7
Starling 7B Zhu et al. (2024a) 7
Jetmoe 8B Shen et al. (2024) 8
GLM 4 9B GLM et al. (2024) 9
Mistral 8B Jiang et al. (2023) 8
@@ 4.3 Implementation Details
    Inference Environment: Most models were accessed via the HuggingFace transformers library using RTX 6000 GPU. DeepSeek 236B model was evaluated using the official API.
@@
    Dataset Rephrasing Prompt: We instruct the rephrasing tool using the following prompt to generate an alternative version of each question while preserving its original meaning and correct answer: “Rephrase the following question without changing its context or the correct answer: {question}”
@@
    Query Prompt: For every query, we construct a standardized input by prepending a fixed instruction to the original MMLU question. Importantly, the multiple-choice options remain identical between the original and the rephrased forms. The fixed instruction is:
    “Select the best answer from the given options. Respond with only the letter corresponding to the correct choice.
    Question: {question}”
@@ 4.4 Evaluation Metrics
We assess model performance by comparing the original dataset, $\mathcal{D}$, with its perturbed counterpart, $\mathcal{D}_{1.0}$, using the following metrics:
@@
Accuracy=#Correct Predictions#Total $\text{Accuracy}=\frac{\#\text{Correct Predictions}}{\#\text{Total Samples}}.$
Absolute and Percentage Performance Difference: The absolute difference in the number of correct answers between $\mathcal{D}$ and $\mathcal{D}_{1.0}$ is denoted by $\Delta_{1.0}$; we also report the relative difference. Statistical Significance: McNemar’s test is applied on the paired predictions to determine whether the performance gap is statistically significant ($p<0.05$)
4.5 Reproducibility
C-BOD source code and datasets, including scripts for data pre-processing, perturbation generation, perturbed datasets, model evaluation, and statistical analysis, are publicly available [^4] https://github.com/SeffiCohen/CBOD. This ensures that our experiments can be independently replicated and verified.
@@ 5 Results
5.1 Overall Performance
As shown in Table 2, most models (20 out of 26) exhibit a noticeable drop in performance on the rephrased test set compared to the original, reinforcing our motivation that these LLMs overfit to the standard MMLU format. Notably, the Llama 1B, Llama 3B models maintain relatively stable accuracy, suggesting they are less susceptible to overfitting. We also observed that Falcon 7B, DeepSeek 7B, Qwen 2.5 3B and Jetmoe 8B show statistically insignificant differences, likely due to their lower baseline accuracy. McNemar’s test confirms that the performance declines observed in most models are statistically significant. Notably, no model shows a significant improvement when inputs are rephrased. This indicates that the C-BOD method reliably uncovers model vulnerabilities rather than occasionally yielding unintentional performance gains.
@@ 5.1 Overall Performance
Across all evaluated models, the average drop in accuracy was 2.15%, and when considering only the models with statistically significant differences, this drop increased to 2.72%.
@@
Table 2: Comparison of LLM performance on the original and perturbed MMLU datasets ($\mu=1.0$). The table shows the number of correct answers on each dataset, accuracy, the absolute and percentage performance difference ($\Delta_{1.0}$), statistical significance ($p<0.05$), and whether the model performed better on the original or perturbed dataset. Models are sorted by parameter count (ascending).
Model Name Model Family Par (B) $\mathcal{D}$ Correct $\mathcal{D}$ Accuracy $\mathcal{D}_{1.0}$ Correct $\mathcal{D}_{1.0}$ Accuracy # $\Delta_{1.0}$ $\Delta_{1.0}$ Sig. Better Perf.
Llama 3.2 1B Llama 1 3799 26.98 3802 27.00 -3 -0.08 No Not Sig
Gemma 2 2B Gemma 2 6656 47.28 6552 46.54 104 1.56 Yes Original
Qwen 2.5 3B Qwen 3 5836 41.45 5739 40.76 97 1.66 No Not Sig
Llama 3.2 3B Llama 3 7940 56.40 7989 56.74 -49 -0.62 No Not Sig
Phi 3.5 4B Phi 4 9547 67.81 9325 66.23 222 2.33 Yes Original
Qwen 2.5 1.5B Qwen 5 5137 36.49 4986 35.41 151 2.94 Yes Original
Yi 1.5 6B Yi 6 8899 63.21 8525 60.55 374 4.20 Yes Original
Qwen 2.5 7B Qwen 7 6990 49.65 6810 48.37 180 2.58 Yes Original
Gemma 7B Gemma 7 8173 58.05 8050 57.18 123 1.50 Yes Original
Apollo2 7B Apollo2 7 9547 67.81 9146 64.96 401 4.20 Yes Original
Aquila 7B Aquila 7 4586 32.57 4475 31.78 111 2.42 Yes Original
Bloomz 7B Bloomz 7 6251 44.40 6133 43.56 118 1.89 Yes Original
Falcon 7B Falcon 7 3733 26.51 3718 26.41 15 0.40 No Not Sig
Starling 7B Starling 7 8364 59.41 8179 58.09 185 2.21 Yes Original
DeepSeek 7B DeepSeek 7 6049 42.96 6077 43.16 -28 -0.46 No Not Sig
Llama 3.1 8B Llama 8 6290 44.68 6169 43.82 121 1.92 Yes Original
Mistral 8B Mistral 8 6928 49.21 6691 47.52 237 3.42 Yes Original
Jetmoe 8B Jetmoe 8 6162 43.77 6132 43.55 30 0.49 No Not Sig
GLM 4 9B GLM 9 9674 68.71 9346 66.38 328 3.39 Yes Original
Yi 9B Yi 9 9345 66.38 9180 65.20 165 1.77 Yes Original
Phi 4 15B Phi 15 10776 76.54 10530 74.79 246 2.28 Yes Original
DeepSeek V2 16B DeepSeek 16 7466 53.03 7358 52.26 108 1.45 Yes Original
Gemma 27B Gemma 27 10300 73.16 9903 70.34 397 3.85 Yes Original
Qwen 2.5 32B Qwen 32 11262 80.00 10816 76.82 446 3.96 Yes Original
Qwen 2.5 72B Qwen 72 11456 81.37 11081 78.71 375 3.27 Yes Original
DeepSeek 236B DeepSeek 236 10648 75.63 10292 73.10 356 3.34 Yes Original
@@ 5.2 Relationship Between Model Size and Overfit Detection
Figure 3 illustrates the scatter plot of the percentage performance difference versus the number of parameters, with a red dashed line representing the logarithmic fit ($\Delta_{1.0}=0.6318\cdot\ln\bigl{(}\text{\# Params}\bigr{)}+0.7920$). The significant log-linear relationship indicates that the performance difference increases with model size in a logarithmic fashion, suggesting diminishing returns as the number of parameters grows.
Figure 3: Scatter plot of the performance difference ($\Delta_{1.0}$) versus the number of model parameters (log scale). A logarithmic trendline is shown. Different colors represent different model families, highlighting how scaling affects robustness to perturbations.
Figure 4 plots the percentage performance difference ($\Delta_{1.0}$) for $\mu=1.0$ against the number of model parameters, with separate plots for models below and above 10B parameters (different x-axis scales). The data reveals a positive trend: larger models tend to exhibit greater performance degradation under textual perturbations. For example, models in the Gemma family show a progressive increase in $\Delta_{1.0}$ with higher parameter counts, while Llama models maintain low $\Delta_{1.0}$ values across scales. The dotted trend line further highlights this relationship.
Figure 4: Performance difference ($\Delta_{1.0}$) versus the number of model parameters for models with (a) fewer than and (b) more than 10 billion parameters. A trendline is shown, and different colors represent different model families, illustrating how model scale within each family relates to the impact of perturbations.
@@ 5.3 Relationship Between Model Accuracy and Overfit Detection
Figure 5 examines the relationship between baseline accuracy on the original prompts and the corresponding percentage difference in performance when evaluated on rephrased inputs. The plot clearly indicates that models with higher original accuracy tend to experience larger declines when exposed to prompt perturbations. For example, a model achieving over 80% accuracy on the original set shows one of the largest $\Delta_{1.0}$ values, while models with lower baseline accuracy exhibit only minor, often statistically insignificant, differences.
This observation highlights a paradox in current LLM evaluation: models that perform exceptionally well on standard benchmarks may be capitalizing on dataset-specific cues rather than demonstrating robust language understanding. The positive correlation between original accuracy and $\Delta_{1.0}$ underscores the need to carefully interpret high benchmark scores, as they might mask underlying vulnerabilities to prompt variations.
@@
Table 3: Examples of how rephrasing affects LLM performance, illustrating potential overfitting to specific phrasing in the original MMLU dataset. The table shows original and rephrased questions, along with an explanation of why the model’s prediction changed. The examples are from Qwen2.5 (32B parameters).
@@ 6.1 Why Do LLMs Overfit?
6.2 Forget What You Know About LLM Evaluation
Ideally, LLMs should exhibit resilience when faced with variations in prompt wording and structure. In other words, robust LLMs are expected to maintain their performance regardless of how a question is phrased, thereby reflecting true language understanding rather than mere memorization. However, our experiments reveal a contrary trend: models that score highly on standard benchmarks often display heightened sensitivity to even minor alterations in prompt formulation. This behavior suggests that such models have implicitly overfitted to the specific linguistic patterns and structures characteristic of these benchmarks. As a result, when these surface-level cues are modified, performance declines, a phenomenon that underscores the paradox between high benchmark accuracy and genuine generalization.
@@ 6.2 Forget What You Know About LLM Evaluation
Although we used MMLU as a demonstration, our approach is inherently dataset-agnostic. It can be applied to any benchmark by simply adapting the performance metric used to compare the original samples with their rephrased counterparts.
@@ 7 Conclusion
By providing a practical, dataset-agnostic framework, our work equips the community with a powerful tool to uncover overfitting and to drive the development of benchmarks that better capture genuine generalization. Incorporating these parametric transformations into the evaluation process not only exposes hidden vulnerabilities in current LLMs but also suggests a way for the creation of more resilient models that can adapt to the evolving challenges of language tasks.
@@ 8 Limitations
Acknowledgements
We used ChatGPT-4o for editing the language and refining the presentation of the text in this paper. The authors affirm that all research content and ideas are their own, and they take full responsibility for the final submitted manuscript.

### v2

@@ 4.2 Models Under Evaluation
Table 4 in Appendix B provides an overview of the LLMs evaluated in our experiments. Our study covers a diverse set of architectures and parameter scales ranging from 1B to 27B parameters. This broad selection enables an in-depth analysis of how both architectural choices and model scale affect robustness to prompt perturbations.
@@ 4.3 Implementation Details
    Inference Environment: All open-weight models were accessed via the HuggingFace transformers library using RTX 6000 GPU.
@@
    Dataset Rephrasing Prompt: We instruct the rephrasing tool using the prompt detailed in Appendix A.1.
@@
    Query Prompt: For every query, we construct a standardized input by prompting a fixed instruction to the original benchmark dataset question. Importantly, the multiple-choice options remain identical between the original and the rephrased forms. The fixed instruction is:
    “Select the best answer from the given options. Respond with only the letter corresponding to the correct choice. Question: {question}”
@@ 4.4 Evaluation Metrics
We assess model performance by comparing the original dataset, $\mathcal{D}$, with its perturbed counterpart, $\mathcal{D}_{0.5}$, $\mathcal{D}_{1.0}$, using the following metrics:
@@
$\text{Accuracy}=\frac{\#\text{Correct Predictions}}{\#\text{Total Samples}}.$
Absolute and Percentage Performance Difference: The absolute difference in the number of correct answers between $\mathcal{D}$ and $\mathcal{D}_{0.5}$ is denoted by $\Delta_{0.5}$; we also report the relative difference. Statistical Significance: McNemar’s test is applied on the paired predictions to determine whether the performance gap is statistically significant ($p<0.05$).
@@ 5 Results
5.1 Open Weight Models - Overall Performance
As shown in Table 1, the majority of models (22 out of 27 for $D_{0.5}$ and 20 out of 27 for $D_{1.0}$) exhibit a noticeable drop in performance on the rephrased test set compared to the original, supporting our hypothesis that many LLMs are sensitive to prompt structure. Notably, smaller models like Llama 1B and Llama 3B maintained relatively stable accuracy, suggesting they are less prone to overfitting, potentially due to their more limited capacity for memorizing superficial patterns.
We also observed that models with lower baseline accuracy tend to show statistically insignificant differences, likely because their initial performance leaves less room for detectable degradation. Importantly, McNemar’s test confirmed that the observed performance drops in most models were statistically significant ($p<0.05$), reinforcing the reliability of our method.
@@ 5.1 Open Weight Models - Overall Performance
Across all evaluated models, the average drop in accuracy for $D_{0.5}$ was 2.15%, which increased to 2.75% when considering only models with statistically significant differences. For $D_{1.0}$, the average drops were 1.87% overall and 2.78% for models with significant performance changes, underscoring the broader impact of stronger perturbations.
@@
Table 1: Comparison of LLM performance on the original and perturbed MMLU datasets. Models are sorted by parameter count (ascending).
Model
Name
Par
$\mathcal{D}$
Accuracy
Claude (0.5) DeepSeek (1.0)
$\mathcal{D}_{0.5}$
Accuracy
$\Delta_{0.5}$
$\Delta_{0.5}$
Better
0.5
$\mathcal{D}_{1.0}$
Accuracy
$\Delta_{1.0}$
$\Delta_{1.0}$
Better
1.0
Gemma-2B 2 47.28 46.47 124 1.86 Original 46.54 104 1.56 Original
Gemma-4B 4 66.73 54.78 222 2.80 Original 65.07 234 2.49 Original
Gemma-7B 7 58.05 55.18 127 1.61 Original 57.18 123 1.50 Original
Gemma-12B 12 68.71 68.77 320 3.20 Original 66.38 328 3.39 Original
Gemma-27B 27 73.16 70.80 356 3.45 Original 70.34 397 3.85 Original
Llama-1B 1 28.11 28.09 3 0.08 Not Sig 27.00 -3 -0.08 Not Sig
Llama-3B 3 56.12 55.49 89 1.13 Not Sig 56.74 -49 -0.62 Not Sig
Llama-8B 8 45.93 45.21 102 1.58 Original 44.68 121 1.92 Original
Mistral-7B 7 57.48 56.33 163 2.01 Original 56.31 128 1.59 Original
Mistral-8B 8 68.05 65.94 298 3.10 Original 66.01 262 2.95 Original
Phi-3.8B 3.8 56.42 56.01 57 0.72 Not Sig 56.31 -44 -0.56 Not Sig
Phi-14.7B 14.7 76.77 74.69 294 2.72 Original 74.79 246 2.28 Original
Phi-14.7B-Reasoning 14.7 73.94 72.05 303 2.90 Original 72.93 142 1.36 Original
Qwen-0.6B 0.6 39.29 39.20 12 0.22 Not Sig 39.48 -33 -0.60 Not Sig
Qwen-1.5B 1.5 38.23 36.46 249 4.63 Original 35.41 151 2.94 Original
Qwen-1.7B 1.7 52.98 51.94 147 1.97 Original 52.20 91 1.22 Not Sig
Qwen-3B 3 41.49 41.46 5 0.09 Not Sig 40.76 97 1.66 Not Sig
Qwen-4B 4 66.97 64.27 381 4.04 Original 65.07 234 2.49 Original
Qwen2.5-7B-Instruct 7.0 58.72 56.92 253 3.06 Original 48.37 180 2.58 Original
Qwen3-8B 8.0 69.73 67.83 267 2.72 Original 67.82 251 2.56 Original
Qwen3-14B 14.0 66.86 65.87 139 1.48 Original 66.11 17 0.18 Not Sig
Yi-6B 6 63.19 60.63 361 4.06 Original 60.55 374 4.20 Original
Yi-9B 9 66.38 65.17 170 1.82 Original 65.20 165 1.77 Original
Apollo-7B 7 67.81 65.42 337 3.53 Original 64.96 401 4.20 Original
GLM-9B 9 68.71 66.06 296 3.06 Original 66.38 328 3.39 Original
Starling-7B 7 59.41 58.15 177 2.12 Original 58.09 185 2.21 Original
Zephyr-7B 7 56.25 54.92 194 2.45 Original 55.05 169 2.13 Original
@@ 5.2 Relationship Between Model Size and Overfit Detection
Figures 3, 4 illustrate the scatter plot of the percentage performance difference versus the number of parameters, with a red dashed line representing the logarithmic fit. The significant log-linear relationship indicates that the performance difference increases with model size in a logarithmic fashion, suggesting diminishing returns as the number of parameters grows. The data reveals a positive trend: larger models tend to exhibit greater performance degradation under textual perturbations. For example, models in the Gemma family show a progressive increase in $\Delta_{1.0}$ with higher parameter counts. The dotted trend line further highlights this relationship.
Figure 3: Scatter plot of the performance difference ($\Delta_{0.5}$) versus the number of model parameters. A logarithmic trendline is shown:
$\Delta_{0.5}=0.6090\cdot\ln\bigl{(}\text{\# Params}\bigr{)}+1.303$.
Figure 4: Scatter plot of the performance difference ($\Delta_{1.0}$) versus the number of model parameters. A logarithmic trendline is shown:
$\Delta_{1.0}=0.7433\cdot\ln\bigl{(}\text{\# Params}\bigr{)}+0.6406$.
@@ 5.3 Relationship Between Model Accuracy and Overfit Detection
Figures 5, 6 examine the relationship between baseline accuracy on the original prompts and the corresponding percentage difference in performance when evaluated on rephrased inputs. The plot clearly indicates that models with higher original accuracy tend to experience larger declines when exposed to prompt perturbations. For example, models achieving over 60% accuracy on the original set present the largest $\Delta_{0.5}$, $\Delta_{1.0}$ values, while models with lower baseline accuracy exhibit only minor, often statistically insignificant, differences.
This observation highlights a paradox in current LLM evaluation: models that perform exceptionally well on standard benchmarks may be capitalizing on dataset-specific cues rather than demonstrating robust language understanding. The positive correlation between original accuracy and $\Delta_{\mu}$ underscores the need to carefully interpret high benchmark scores, as they might mask underlying vulnerabilities to prompt variations.
Figure 5: Scatter plot showing $\Delta_{0.5}$ for $\mu=0.5$ against the original accuracy of the model. Models within the same family are marked with the same color.
@@
Table 2: Examples of how rephrasing affects LLM performance, illustrating potential overfitting to specific phrasing in the original MMLU dataset. The table shows original and rephrased questions, along with an explanation of why the model’s prediction changed. The examples are from Qwen3 (14B parameters).
@@
High-School Biology “Which of the following is NOT a characteristic of bacteria?” “Which of the listed options fails to represent a defining trait of bacterial organisms?” The explicit cue word “NOT” is replaced by the longer clause “fails to represent.” LLM appears to rely on a pattern like “Which … is not …” to flip polarity; when the negator is hidden inside a relative clause, the learnt template does not fire, so the model picks a true trait instead of the exception.
High-School Mathematics “Positive integers $x,y$ satisfy $xy=56$, $x<y$, and $7$ divided by the reciprocal of the larger integer equals $4$. What is xx?” $x,y$ are positive integers with product 56 and $x<y$. If seven divided by the larger integer results in 4, determine xx.” The word “reciprocal” pins a template: $7\times\frac{1}{y}=4\Rightarrow y=\frac{7}{4}$. Replacing it with a looser “seven divided by the larger integer results in 4” flips the parse—The LLM treats “results in” like a remainder cue (7divy=47\ $\text${div}\ y=4), producing $y=1$ and thus the wrong $x$.
@@
5.4 Closed-Source Models: Accuracy vs. Robustness Trends
We analyze the performance of proprietary GPT models (GPT-4o [^2] https://openai.com/index/gpt-4o-system-card/?utm_source=chatgpt.com, GPT-4.1333https://openai.com/index/gpt-4-1/?utm_source=chatgpt.com, GPT-5444https://openai.com/index/gpt-5-system-card/?utm_source=chatgpt.com) under the C-BOD rephrasing evaluation. While these models achieve high accuracy on unperturbed MMLU prompts, they still show significant degradation under semantically equivalent prompt modifications ($\mu=0.5$), as shown in Table 3.
Model Accuracy $\Delta_{0.5}$ $\Delta_{0.5}$ Better
$\mathcal{D}$ $\mathcal{D}_{0.5}$
GPT-4o 83.88% 81.46% 339 2.88 Original*
GPT-4o-mini 74.96% 72.37% 364 3.46 Original*
GPT-4.1-nano 69.68% 67.53% 302 3.09 Original*
GPT-5-nano 86.63% 84.30% 326 2.68 Original*
GPT-5-mini 90.56% 88.35% 310 2.44 Original*
Table 3: Performance of closed-source GPT models on MMLU under prompt perturbation ($\mu=0.5$). All results are statistically significant (marked with *) according to McNemar’s test (p < 0.05).
Figure 7: Accuracy vs. Drop for GPT models. Each point shows original accuracy and percentage drop after rephrasing.
Closed-source models are sensitive to rephrasing.
Despite their high original accuracy, all closed-source GPT models demonstrate significant performance degradation under prompt rephrasing. The observed drops range from 2.44% (GPT-5-mini) to 3.46% (GPT-4o-mini). This reinforces the hypothesis that these models exploit benchmark-specific surface patterns.
Newer models perform better.
A clear upward trend is observed across model generations: GPT-5-mini achieves the highest original accuracy (90.56%) and also shows relatively lower sensitivity (2.44% drop) compared to older GPT-4o models (drops of 2.88–3.46%). This suggests that newer training pipelines or architectures improve generalization.
Larger models generalize better.
Within each family, the larger variants ("mini") outperform their smaller counterparts ("nano") both in raw accuracy and robustness. For instance, GPT-5-mini achieves +3.93% higher accuracy than GPT-5-nano (90.56% vs. 86.63%) while maintaining a similar drop (2.44% vs. 2.68%), indicating effective scaling without sacrificing generalization. Likewise, GPT-4o-mini outperforms GPT-4.1-nano by +5.28% accuracy while showing only a slightly higher sensitivity to rephrasing.
@@ 6 Discussion
Forget What You Know About LLMs Evaluation
Ideally, LLMs should exhibit resilience when faced with variations in prompt language and structure. In other words, robust LLMs are expected to maintain their performance regardless of how a question is phrased, thereby reflecting true language understanding. However, our experiments reveal a contrary trend: models that score highly on standard benchmarks often display heightened sensitivity to even minor alterations in prompt formulation. This behavior suggests that such models have implicitly overfitted to the specific linguistic patterns and structures of these benchmarks. As a result, when these surface-level cues are modified, performance declines, a phenomenon that underscores the paradox between high benchmark accuracy and genuine generalization.
@@
Although we used MMLU and GPQA as a demonstration, our approach is inherently dataset-agnostic. It can be applied to any benchmark by simply adapting the performance metric used to compare the original samples with their rephrased counterparts.
@@ 7 Conclusion
By offering a practical, dataset-agnostic framework, this work equips the community with a robust tool to identify overfitting and foster the development of benchmarks that more effectively assess genuine generalization. Integrating these parametric transformations into the evaluation process reveals hidden vulnerabilities in existing LLMs and paves the way for designing more resilient models capable of adapting to the ever-evolving challenges of language tasks.
@@ 8 Limitations
Appendices
Appendix A Prompts Used in CBOD Framework
A.1 Distortion Prompt
The rephrasing prompt was developed using a meta-prompting approach, which involves iteratively refining the prompt based on systematic error analysis. This process aimed to progressively improve the quality of the generated perturbations, ensuring high semantic fidelity and consistency. For example, early evaluations revealed approximately 25 cases where prompts involving sensitive content (e.g., sexual topics) led to inaccurate or off-target rephrasings. In response, we refined the prompt to explicitly exclude such cases, effectively addressing this error class and improving overall performance. This iterative tuning, incorporating both automated and human feedback, ultimately resulted in 100% accuracy as measured by the reasoning model judge. The prompt:
system: "You are a rephrasing assistant tasked with preserving the original structure, type, and intent of technical questions or statements. Your goal is to rephrase while maintaining meaning, purpose, structure, and context. keep the same level of clarity, do NOT try to simplify. keep question a question / statement a statement. do NOT try to answer the question" user: "Rephrase the following question in different wording, ensuring the meaning remains exactly the same. Match the readability level. Do NOT change the type of sentence: if it is a definition, keep it a definition; if it is a multiple-choice question, keep it as such; if it is a procedure or description, keep it in that form. Do NOT infer, guess, or introduce new information, assumptions, or constraints. Do NOT reword a description into a question, or vice versa. Keep all scientific and technical terms, units, variables, equations, and formatting intact. Your output should differ in phrasing only, not in meaning. Do NOT improve clarity or readability. Do NOT try to answer the question. Sexual or harmful content rephrasing is allowed for academic purpose. Return ONLY the rephrased version with no extra text or commentary. TEXT to rephrase: {question}"
A.2 Judge Prompt
You will be given two versions of a question: an ‘original_question‘ and a ‘rephrased_question‘. Your task is to evaluate if they have the exact same meaning. "Semantically equivalent" means that a person with the required domain knowledge would provide the exact same answer to both questions. Respond with a single JSON object containing two keys: 1. ‘"judgment"‘: Your verdict, which must be either "EQUIVALENT" or "NOT_EQUIVALENT". 2. ‘"reasoning"‘: A brief, one-sentence explanation for your judgment. Judgment Criteria: A rephrased question is NOT_EQUIVALENT if it meets any of the following conditions: A. Logical Alteration * Reverses Logic: The rephrase swaps the subject and object or reverses the direction of an implication. * Changes Logical Operator**: The rephrase changes a one-way implication (if/then) to a two-way bi-conditional (if and only if). B. Change in Scope or Precision * Loss of Specificity: The rephrase replaces a precise technical term with a vague or overly general one. * Incorrect Substitution: The rephrase swaps a key term with another, related but incorrect, term (e.g., "mass" for "weight"). * Incorrect Expansion of Acronym: The rephrase incorrectly defines an acronym for the given context. C. Structural Failure * Answers the Question: The rephrase provides the definition or answer to the original question, especially in fill-in-the-blank scenarios. * Fails to Rephrase: The output is an error message, a refusal, or otherwise not a good-faith attempt at rephrasing. A rephrased question is EQUIVALENT only if it avoids all the errors above. Stylistic changes, synonym swaps, and sentence restructuring are acceptable as long as the core meaning remains identical.
Appendix B Models Tested
Table 4 presents all the LLMs we tested with their versions.
Table 4: Overview of the evaluated open-weight LLMs. Models are grouped by family, model version, and the number of parameters (in billions).
Family Version Params
Qwen Qwen2.5 1.5B Yang et al. (2024) 1.5
Qwen2.5 3B 3
Qwen2.5 7B 7
Qwen3 0.6B 0.6
Qwen3 1.7B 1.7
Qwen3 4B 4
Qwen3 8B 8
Qwen3 14B 14
Llama 3 Llama 3.2 1B Dubey et al. (2024) 1
Llama 3.2 3B 3
Llama 3.1 8B 8
Gemma Gemma 2 2B Team et al. (2024) 2
Gemma 4B 4
Gemma 7B 7
Gemma 12B 12
Gemma 27B 27
Phi Phi 4 4B Abdin et al. (2024) 4
Phi 4 15B 15
Phi 4-reasoning 15B 15
Mistral Mistral 7B Jiang et al. (2023) 7
Mistral 8B 8
Yi Yi 6B Young et al. (2024) 6
Yi 9B 9
Others Apollo2 7B Zhu et al. (2024b) 7
Starling 7B Zhu et al. (2024a) 7
GLM 4 9B GLM et al. (2024) 9
Zephyr 7B Tunstall et al. (2023) 7
Appendix C Additional benchmark - GPQA results
Table 5: Comparison of LLM performance on the original and perturbed GPQA datasets. Models are sorted by parameter count (ascending). We report different levels of $\mu$.
Model
Name
Par
$\mathcal{D}$
Accuracy
$\mathcal{D}_{0.5}$
Accuracy
$\Delta_{0.5}$
Better
0.5
$\mathcal{D}_{1.0}$
Accuracy
$\Delta_{1.0}$
Better
1.0
$\mathcal{D}_{1.5}$
Accuracy
$\Delta_{1.5}$
Better
1.5
Gemma 2B 2 34.25 29.12 14.93 Original 30.95 9.60 Original 29.30 14.46 Original
Gemma 7B 7 46.52 41.58 10.62 Original 41.94 9.83 Original 38.64 16.91 Original
Llama-3B 3 36.45 33.15 9.05 Original 33.88 7.04 Not Sig 32.23 11.56 Original
Llama-3B-Instruct 3 21.06 21.43 1.76 Not Sig 17.58 16.51 Original 20.15 4.32 Not Sig
Phi-4-reasoning-plus 14.7 16.85 15.02 10.86 Original 15.30 9.07 Original 15.02 10.86 Original
Qwen 4B 4 56.04 53.84 3.92 Not Sig 52.56 6.20 Original 53.29 4.90 Not Sig
To assess the robustness and generalizability of the proposed C-BOD method, we conducted an ablation study on the GPQA benchmark using varying levels of textual distortion ($\mu$). Table 5 reports the performance of leading LLMs on both the original and rephrased GPQA datasets, where rephrasings were generated by GPT-4o with $\mu\in\{0.5,1.0,1.5\}$.
The results on GPQA are mostly consistent with those observed on MMLU. Across multiple LLM families and distortion levels, we observe a systematic performance drop on the rephrased inputs, suggesting a sensitivity to surface-level prompt variations. While the degree of statistical significance varies by model and $\mu$, the overall trend of declining accuracy under rephrasing reinforces the hypothesis that many models rely heavily on benchmark-specific prompt formulations.
At the same time, several characteristics of GPQA help explain why the results differ from those obtained on MMLU and why it is more difficult to show consistent effects across all models at this stage. First, GPQA is a very new benchmark, specifically designed for reasoning, and thus has not yet been exposed widely enough to influence model training. Second, most models achieve very low absolute accuracy on GPQA, often failing to surpass 20%. At such a low baseline, small fluctuations in predictions can obscure clear differences and make it harder to observe statistically significant effects. Third, GPQA is a relatively small dataset, which further limits the statistical power of tests such as McNemar’s and reduces our ability to detect significance robustly.
Overall, whenever some of the models keep the expected trend, some models suffer from very low accuracy. This, combined with the novelty of the benchmark, its reasoning-oriented design, and its small size, explains why it is currently more difficult to show across-the-board results. Importantly, the uniformly low accuracy also suggests that overfitting is not yet strongly affecting GPQA, as models have not learned enough benchmark-specific artifacts to inflate their scores in the first place. Instead, their failures likely stem from the underlying reasoning challenges posed by the dataset, making GPQA a valuable complement to more saturated benchmarks like MMLU.